# Understanding the Skill Gap in Recurrent Language Models: The Role of the Gather-and-Aggregate Mechanism

### v1

@@ A.5 Retrieval Heads Across Model Families
Table 7: Attention Head Ablation Study: Performance impact on KV-Retrieval task when individual attention heads are removed from various language models. Each head was systematically knocked out and the model’s accuracy was evaluated on a KV-Retrieval benchmark with dictionary size of 20 across 1000 samples. After measurement, each head was restored before testing the next one. Results show the 50 most significant attention heads from each model, sorted by accuracy, revealing which heads contribute most critically to the KV-Retrieval capability across Zamba2, Llama, and Llamba model families. For Zamba2 models, marks layers with the first shared attention, and marks those with the second.

### v2

@@ A.6 Retrieval Heads Across Model Families
Table 5: Attention Head Ablation Study: Performance impact on KV-Retrieval task when individual attention heads are removed from various language models. Each head was systematically knocked out and the model’s accuracy was evaluated on a KV-Retrieval benchmark with dictionary size of 20 across 1000 samples. All measurements were performed using ‘Answer scoring’ without a trailing space in the prompt. After measurement, each head was restored before testing the next one. Results show the 50 most significant attention heads from each model, sorted by accuracy, revealing which heads contribute most critically to the KV-Retrieval capability across Zamba2, Llama, and Llamba model families. For Zamba2 models, marks layers with the first shared attention, and marks those with the second.